Instrument Log

 

NIMFR C1        Central Facility, Lamont, OK

 
 

Date

Technician

Comments

11/10/15

JEM

BCR-02224 requests that the NIMFR be removed and sent to Boulder.  Removed the *NIMFR sensor, control box and cabling.  Disabled NIMFR collections at 1945 GMT.  Took the instrument to shipping and receiving to be shipping to Gary Hodges at Boulder.

*Old S/N:043 WD20276

2/11/15

JEM

Mentor indicates the temperature is not being regulated properly and is getting too high.  R/R the *loggerboard as well as the tube temperature controller.  RTS @1830 GMT.   *Old - $3A0A   *New - $9E12

2/6/15

JEM

The NIMFR needs to be installed.  Installed the *NIMFR and **loggerboard @2030 GMT.  Initialized the NIMFR and verified that all channels were reading.  RTS @2100 GMT. 

*New - $F799   **New - $3A0A

10/1/14

JEM

Due to continued problems with the NIMFR it has been removed from service until bench testing and repairs can be completed and data is accurate.

9/26/14

JEM

The site had a temporary power outage for some electrical working.  The power outage cause an hour of missing data between 1300 GMT and 1400 GMT.  RTS @1400 GMT.

9/24/14

JEM

The site had a temporary power outage for some electrical working.  The power outage cause an hour of missing data between 1400 GMT and 1500 GMT.  RTS @1500 GMT.

9/22/14

JEM

The NIMFR channel 940 is still over-ranging and all others are too high.  Removed the NIMFR sensor for testing on the Standard Lamp Bench.  Testing found that the sensor itself is what is causing the high readings.  Repaired sensor head and verified all readings were in range on the Standard Lamp Bench before reinstalling the sensor.  Reinstalled the sensor and restarted collections @2100 GMT.

9/10/14

JEM

All lab tests for board 3A0A were OK, but in the meantime the original board (B074) has been repaired and will be reinstalled.  Installed the *logger baord and restarted collections @2130 GMT.     *New - B074

9/8/14

JEM

DQPR #4278 indicates that the current logger board may have the incorrect gain resistors installed because channel 940 is over-ranging and all of the others are too high.  Removed the *logger board @1300 GMT to test and verify gain resistors in the lab.  The NIMFR will remain down while the board is tested. Did not RTS.  *Old - 3A0A

8/8/14

JEM

While checking the NIMFR remotely I noticed that the detector temperature was maxed out at 4096 counts.  Troubleshooting found that the leads for the detector thermistor are shorted together.  Further troubleshooting determined that the problem was inside the actual sensor.  Removed the sensor to continue troubleshooting in the lab.  Determined that the thermistor itself had failed.  R/R the thermistor and tested.  Reassembled the sensor and reinstalled it at 1730 GMT.  Both temperatures are reading properly now.  RTS @1740 GMT.

8/8/14

JEM

The NIMFR began failing collections on 8/7/14 at 2008 GMT.  The logger board, SN - B074, has failed again. R/R the *logger board.  The new logger board is working, but the detector temperature is maxed out at 4096 counts.  This indicated a failed thermistor or a short in the 2 leads for the thermistor somewhere.  Determined that the problem was in the sensor somewhere and not in the sensor cabling.  Removed the NIMFR sensor and returned to the lab for further troubleshooting.  Troubleshooting found that the thermistor itself had failed.  R/R the thermistor and reassembled the sensor.  Reinstalled the sensor @1445 GMT.  Initialized the NIMFR @1453 GMT.  RTS @1500 GMT.

*Old – B074  *New – 3A0A

8/5/14

JEM

The tube and detector temperatures are dropping to 0 C as indicated in DQPR #4255.  It seems as if there is a short in the board because the temperatures drop to 0 C, but then the temperatures also closely trend with the detectors.  Logger board SN - B074 was repaired and can be reinstalled.  Installed the new *logger board at 2120 GMT.  Notified the mentor and DMF folks of the change.  RTS @2130 GMT.   *New – B074

7/28/14

JEM

The NIMFR has still not collected since it was shut down for the SIRS upgrade.  Further troubleshooting determined that the *data logger needs to be replaced.  Made modifications to a spare datalogger to enable it to be connected to a new style head like the one in the current NIMFR.  Installed the modified datalogger and communications tested OK.  Data is now being collected.  Notified the mentor of the new logger board so updated calibration files can be generated.  RTS @1800 GMT.

*Old - B074    *New - B000

7/25/14

JEM

While the SIRS ventilators are upgraded the NIMFR power needs to be disconnected.  Disconnected NIMFR power at 1400 GMT.  NIMFR power was restored at 1809 GMT. Unable to establish communications with the NIMFR logger.  Trouble shooting determined that the fiber optics were not damaged during the upgrade.  Also communications cannot be established using a laptop connected directly to the datalogger.  Due to time troubleshooting will continue Monday.  RTS @2030 GMT.

6/19/14

JEM

Since the NIMFR maintenance yesterday the tube heater has not been functioning.  Removed and disassembled the NIMFR and found the small 7 conductor Molex connector under the insulation had become disconnected.  Reconnected the Molex connector and taped to prevent future accidental disconnects.  Reassembled the NIMFR and reconnected.  Verified readings and RTS @ 2145 GMT.

6/18/14

JEM

Kenny reported seeing condensation in the NIMFR lens on 6/17/14.  Removed the NIMFR and disassembled it.  Added teflon tape to the threads at each end of NIMFR tube.  After reconnecting the diffuser, flushed the tube with dry air before attaching the lens.  Reassembled the NIMFR and reinstalled it @1505 GMT.  RTS @1515 GMT.

10/8/13

JEM

A communications upgrade needs to be done at E13 and C1.  During this upgrade power was disconnected so that power cords could be rerouted.  NIMFR data is missing from approximately 1700 GMT to 1900 GMT.

8/26/13

JEM

After the logger was replaced and cal set updated with the DMF data ingest resumed.  It was then observed that the head temp was not reading.  Powered the NIMFR off @1508 GMT.  Troubleshooting found several problems.  Two bad solder joints were found on the heater circuit.  A required wire on the sensor cable was cut on the logger board end and needed to be reconnected.  Tried the logger board again but now the instrument will not even establish communication.  Further troubleshooting found a slight short on the logger board where modifications had been made.  2 pins (17 and 16) were supposed to be isolated from each other.  There was a short (1.3k Ω) where some of the trace wasn't cut good.  Completed the isolation of these 2 pins and reinstalled the board.  The board powered up and I initialized the NIMFR.  All readings are present and look normal after checking processed data.  RTS @1825 GMT.

8/19/13

JEM

A new sensor was installed on 8/16/13.  This sensor will not work with the current logger board without modifications being made to it.  R/R the *logger board with one that already has the modifications done.  Powered the system on and initialized the NIMFR @2055 GMT.  RTS @2100 GMT.   *Old - $B432  *New -  $B074

8/16/13

JEM

Mentor requests the NIMFR sensor be replaced with a newly built one.  Powered the system off @1520 GMT. R/R the *NIMFR sensor. Powered the system back on @1655 GMT. Notified mentor that the request had been completed. RTS @1655 GMT.  Old - $5A1F  New - $E984

5/28-29/13

JEM

Instrument mentor wants the NIMFR removed for a quick refurbishment/inspection and to run the head through the standard lamp and spectral benches.  Removed the NIMFR from service to inspect/repair/calibrate.  The NIMFR was reinstalled on 5/29/13 at 1945 GMT after all work was completed.

4/22/13

KT

The entire NIMFR window was covered in bird droppings. At 1550 GMT cleaned NIMFR window. RTS at 1605 GMT.

12/1/09

MK

The NIMFR has been experiencing many problems due to faulty conxall head connector. Now it appears the tube heater has stopped working. After speaking with Gary Hodges is has been decided to remove NIMFR and replace conxall connector and troubleshoot heater problem. Powered off NIMFR on 12/1/09 at 2030 GMT. Disassembled NIMFR sensor in ERL.  R/R conxall connector and wiring harness connecting the cube assembly to the conxall connector. Also found corroded and broken wire leading to one of two tube heaters. Replaced tube heater and rewired heater wires. Resealed, and reinsulated NIMFR sensor and assembly. Reinstalled sensor on 12/3/09 at 2052 GMT.  

12/2/08

CAW

Instrument mentor requested an experiment be carried out which consisted of placing different filters and caps over the MFR sensor for a set time.  Experiment carried out per mentor’s instructions from 1804 to 1815 GMT.  Notified mentor of times for each condition. 

10/11/07

MK

The NIMFR has been repaired and needs to be reinstalled.  Found no voltage getting to pin 2 of the op amp.  No voltage coming through the thermistor.  Found a broken wire inside the insulation on the heater cable. Repaired cable and reinstalled at 1500 GMT, powered on at 1630 GMT.

10/5/07

MK

The heater is not shutting off.  Removed the NIMFR at 1500 GMT for troubleshooting in the electronics lab.

10/5/07

MK

The heater is not shutting off.  Removed the NIMFR at 1500 GMT for troubleshooting in the electronics lab.

6/8/07

MK

The heater is not working at this time. R/R heater fuse at 2038 GMT. RTS at 2045 GMT.

3/20/07

PAD

Power bumps caused by area storms tripped media converters located in Optical trailer and and instrument comms box.  Following instruments were affected:  EBBR, MFRSR, NIMFR, SIRS, SMOS, and SWATS.  Reset media converters located in Optical trailer and at instrument comm box.  RTS at 1953 GMT.

10/31/06

MK

The tube temperature was excessive during the summer months and was removed. Due to the colder weather the fuse should be replaced for further testing. Replaced 3 amp fuse on the tube circuit board. RTS at 2015 GMT.

7/11/06

MK

The NIMFR has stopped communicating with SDS. Checked NIMFR with laptop, instrument is still running normal. Checked fibers from instrument to CC2, checked ok. Checked fibers from CC2 to optical trailer, no response. Found bad port on the fiber splice box. Moved one fiber to different fiber splice, all communications were restored.

6/27/06

MK

The tube temperature dropped to zero on 6/12/06 at 1000 GMT. Found the heater plug disconnected inside the NIMFER sensor enclosure. Connected plug at 1450 GMT. Suspect that at one point the NIMFR cable has hung up on the corner of the Tracker post and pulled the connecter apart. Wired tied all cables together.

6/26/06

MK

The tube temperature dropped to zero on 6/12/06 at 1000 GMT. Checked all board voltages and connectors on data logger board, no problems found.

5/11/06

MK

The NIMFR is back from repairs and needs to be reinstalled. Installed NIMFR at 1815 GMT, powered on at 1830 GMT. Monitored NIMFR, all readings were normal.

4/24/06

MK

Filter detector number 5 673nm is reporting zero volts. Checked all connections on the terminal block and board connector. Channel 5 is reporting 4095. Removed cable from sensor checked plug connection, OK. Plugged the LAB spare MFR head into the NIMFR board, all readings were reporting values. Problem is in the NIMFR sensor head. On 4/24/06 2010 GMT powered off NIMFR logger board and removed NIMFR sensor to be sent in for repair.

3/21/06

PAD

At 1601 GMT, found NIMFR without power.  Power cycled UPS at 1606 GMT.  RTS at 1621 GMT.

1/25/06

MK

The NIMFR has arrived from repair and calibration. Installed NIMFR system at 1930 GMT.  *Logger S/N is *B432.  **Head S/N is **$5A1F.

9/23/05

MR

The instrument mentor wants the system removed from service and shipped back to him. Powered down the system at 1905 GMT.  Removed the system board, head and wiring.  Will ship the system back to the mentor.

8/17/05

MR

The communications between the CF instruments and the SDS has stopped. T/S found that the fiber optic media converter had not reset.  Reset the unit in the optical trailer and in the communications box at E13.  The communications was restored.

8/17/05

MR

DQPR # 583 states that the data is dropping out between 1750 and 2000 GMT daily.  Checked the cables on the instrument and they checked OK.  They did not look kinked or stretched.  Reseated the connections.  RTS at 1805 GMT.

7/6/05

PAD

At 1616 GMT, observed small amount of condensation inside NIMFR lens.  At 1628 GMT, removed lens and cleaned condensation.  RTS at 1643 GMT.

1/18/05

MR

DQPR # 421 requests that the 5 volt head voltage be checked.  Checked the +/- 5 volts to the head and it was missing.  Reseated the head connector and the =/- 5 volts returned.  RTS at 1515 GMT.

7/26/04

PAD

Last successful collection was on 7/23/04 at 2108 GMT.  At 1400 GMT, NIMFR was found without power.  Storms in area on 7/23/04.  Power cycled UPS at 1408 GMT.  RTS at 1423 GMT.

6/10/04

CAM

SDS had not received any data from the NIMFR since 0208 GMT today.  The following error message appears in the collection log:

interact() => Warning: RSR in Unknown state : output -> 

interact() => Warning: Autologin failed - retrying.

This was the same type of error that was occurring prior to power cycling the Rocket Port hub in the control trailer two days ago.  SDS personnel replaced the Rocket Port hub at 1530 GMT.  Forced a collection from the NIMFR at 1536 GMT, which was successful.  RTS status O.K. 

6/8/04

CAM

SDS has not been able to collect data from the NIMFR since 1408 GMT on 6/4.   The following error message appears in the collection log: interact.c:363 => Fatal: (passwd failed) unexpected output from RSR at E13.  The datalogger was replaced on 6/7 with no change in status.  Checked comms from the instrument all the way back to the Control Trailer and was able to connect to and receive a hello prompt from the datalogger with no problems.  The only component remaining that is unique to SDS collections is the Rocket Port hub.  Power cycled the Rocket Port hub in the control trailer at 1925 GMT.  Forced a collection, which was successful at 1929 GMT.  Instrument was successfully restarted and RTS status O.K.

4/20/04

MR

The NIMFR is still unable to communicate with the SDS.  T/S found that the data logger, fiber optics and bit drivers check OK.  Power cycled the rocket port and the communications was restored.  RTS at 2020 GMT.

4/19/04

MR

The NIMFR is unable to communicate with the SDS system.  Power cycled the data logger.  Will wait until next collection to confirm if this repaired the problem.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

10/25/02

Mentor

Mentor reinstalled instrument and RTS at 2000 GMT.

1/9/02

MR

The instrument mentor wants the NIMFR sensor head to be removed and sent back for repairs.  Powered down the instrument at 1445 GMT.  Removed the *sensor head.  Shut the collections and ingest off for this instrument.  Will ship the sensor to John Schmelzer for repairs.  *Old – 5A1FCC

12/10/01

MR

The instrument mentor requests that the data logger be changed out again due to another voltage problem, this time with the instrument head. Changed out the *data logger at 1600 GMT.  Disabled the ingest process for this instrument and notified the mentor of the change.  RTS at 1615 GMT.  *Old - BB2F  *New - 93902

12/6/01

MR

The SDS still is having trouble communicating with the NIMFR.  T/S found that the NIMFR would not communicate with a laptop computer.  Power cycled the instrument at 1533 GMT.  Communications was restored.  SDS was able to communicate with the instrument at 1608 GMT.  RTS at 1615 GMT.

12/5/01

MR

The instrument has not collected any data since 11/29/01.  On 11/30/00 T/S found that the data logger was dead and would not communicate to the SDS or a laptop computer.  Removed data logger and further T/S was performed in the lab.  T/S found that the power supply on the board had stopped working and that further T/S would be needed.  Contacted the mentor on 12/5/01 and he said a spare MFR data logger can be used.  R/R *data logger.  Powered on the instrument at 2130 GMT.   *Old - BB2C   *New - BB2F

11/29/01

MR

The NIMFR tube does not seem to be warm.  The tube should be running at about 40 c.  This indicates that the tube heaters are not working.  Checked the voltage on the heater control board and it checked OK.  Checked the connector on the heater cable and found 3 pins had corrosion on them.  Shut the power off at 1615.  Scrubbed the pins and the socket with a small wire brush to remove the corrosion.  Turned the power back on at 1620.  After the power was restored the tube started to heat.  After about 30 min the tube was checked and it was noticeably warmer than the outside and surrounding metal, as it should be.  RTS at 1645 GMT.

1/22/01

DN

NIMFR power was turned off from 1843 to 1934 GMT to facilitate BSRN/BRS conversion, as the two systems shared the same AC supply.

12/19/00

MR

Instrument mentor is onsite to perform T/S on the NIMFR tube heater circuit. T/S on the tube heater showed nothing wrong with the electronics. Cleaned and reseated the connector and the tube heater began to work better but was not working at maximum efficiency. Mentor suspects that the connections may not be able to handle the current load and may be degrading.

12/12/00

CAM

Due to a hard drive failure on the SDS Development system, data stopped being collected from the NIMFR at 2152 GMT on 12/5/00. The hard drive was replaced and normal data collections resumed at 1840 GMT on 12/12/00 after all the necessary configurations were set.

 

 

Instrument installed by mentor.